OpenStudy (anonymous):

can someone show me how to do this Equation x/3 - 5=y

OpenStudy (anonymous):

When you do basic algebra, the concept is one side must equal the other side of the '=' sign. In order to do this, we must do the same operation to both sides in order to keep the equality. In your example you can see that if you divide both sides by 3 you will get: x/y = 3/5 x / (3*y) = 3 / (3*5) x / (3*y) = 1 / 5 From here we can multiply both sides by 'y' to isolate the x/3: x / (3*y) = 1 / 5 (x*y) / (3*y) = (y*1) / 5 x / 3 = y / 5 So you can now see that your answer is: y / 5 I hope this helps.